# Parity Nonconservation in Rb and Sr+ due to Low-Mass Vector Boson

### v2

@@ I Introduction
The present calculations are performed with accuracy about 1% for 5s6s5s-6s and a few per cent for $5s-4d_{3/2}$ transitions for the Standard model results. We do aim for a higher accuracy here, since the main purpose of the present paper is to determine the relative contribution of new physics effects. All major contributions, such as electron correlation effects, core polarisation, and the Breit interaction are included. At this stage, we neglect smaller corrections, such as QED effects, structure radiation, and the neutron skin contribution.
@@ II PNC amplitudes
Here NN is the number of neutrons, and $Z$ is the number of protons.
@@ III Results and discussion
In both expressions, NN is the neutron number, AA is the mass number, $A_{\rm PNC}$ is the tabulated numerical value of the spin-independent PNC amplitude expressed in units of $iea_{B}$, see Table 1; $A_{\mathrm{PNC}}^{Z^{\prime}}$ is the tabulated numerical value of the $Z^{\prime}$-boson contribution to the PNC amplitude. The values of $A_{\mathrm{PNC}}^{Z^{\prime}}$ used in Eq. (6) should be taken from the first row of Table 3, while those for Eq. (7) correspond to the last row of the table. Parameter $\varepsilon$ denotes the relative deviation of the experimental PNC amplitude from the Standard Model prediction.

### v3

@@ I Introduction
Low sensitivity to neutron skin in light elements is especially important for the isotope chain method suggested in Ref. [13]. In the ratio of PNC effects in different isotopes an error of atomic calculations cancels out. The dominating theoretical error comes from the uncertainties in the neutron distributions. Sensitivity to the neutron skin, investigated in Refs. [14, 15, 16], is proportional to $Z^{2}\alpha^{2}$ and is suppressed in light elements. First experimental implementation in Yb isotope chain was published in Refs. [17, 18]. Strontium has 4 stable isotopes, 84Sr, 86Sr, 87Sr, and 88Sr, and significantly weaker sensitivity to the neutron skin than Yb. Combination of single-isotope and isotope chain methods allows one to measure separately both proton and neutron weak charges.
@@
The present calculations are performed with accuracy about 1% for 5s6s5s-6s and a few per cent for $5s-4d_{3/2}$ transitions for the Standard model results. We do not aim for a higher accuracy here, since the main purpose of the present paper is to determine the relative contribution of new physics effects. All major contributions, such as electron correlation effects, core polarisation, and the Breit interaction are included. At this stage, we neglect smaller corrections, such as QED effects, structure radiation, and the neutron skin contribution.
@@ II PNC amplitudes
Here $N$ is the number of neutrons, and $Z$ is the number of protons.
@@ III Results and discussion
In both expressions, $N$ is the neutron number, $A$ is the mass number, $A_{\rm PNC}$ is the tabulated numerical value of the spin-independent PNC amplitude expressed in units of $iea_{B}$, see Table 1; $A_{\mathrm{PNC}}^{Z^{\prime}}$ is the tabulated numerical value of the $Z^{\prime}$-boson contribution to the PNC amplitude. The values of $A_{\mathrm{PNC}}^{Z^{\prime}}$ used in Eq. (6) should be taken from the first row of Table 3, while those for Eq. (7) correspond to the last row of the table. Parameter $\varepsilon$ denotes the relative deviation of the experimental PNC amplitude from the Standard Model prediction.